OpenStudy (anonymous):

Use set theory or vector notation to describe the points that lie on the given configuration:: The line passing through ( 0,2,1 ) in the direction of 2i-k.

OpenStudy (anonymous):

is it l(t) = (2t)i + 2j + (1-t)k ???

OpenStudy (phi):

Here are some notes http://tutorial.math.lamar.edu/Classes/CalcIII/EqnsOfLines.aspx

OpenStudy (phi):

I think like this: P (for point on the line) is P0 (starting point) + k* direction P= P0 + k D here P0 is ( 0,2,1 ) the direction vector D is (2,0,-1) so P= (0,2,1) + k(2,0,-1)

OpenStudy (phi):

you can use i , j , k instead: P = 0*i + 2*j +1*k + t*2i + t*0j + t*-1k (I switched to using t instead of k, because is now a basis. up above k was just some number) simplify to get P= 2t i + 2j + (1-t) k

OpenStudy (anonymous):

excellent, so i got it correct

OpenStudy (anonymous):

in your first message ( p= 2ti....) what form/notation would you call that?

OpenStudy (phi):

I always find it easier to remember the vector way: start point + t*direction vector you always need a starting point, and it makes sense to add in a direction that you move along however far (using t)

OpenStudy (anonymous):

thank you kindly

OpenStudy (phi):

Paul's notes show how to write this in parametric form or as symmetric equations of a line